OpenStudy (anonymous):

Which of the following equations is equivalent to (5 + x) + 2 = 7? A. 5 + 2x = 7 B. 7x = 7 C. x + 7 = 7 D. 5x + 2 = 7

OpenStudy (mathstudent55):

Drop the parentheses on the left side and add 5 and 2 together.

OpenStudy (anonymous):

(5+x)+2=7 (5+x)=7-2 (5+x)=5 x=3 NONE of the above... when you solve them, x=1, in comparison to your equation x=3 :/

OpenStudy (mathstudent55):

One of the choices is definitely correct.

OpenStudy (anonymous):

yes, unless you drop the parenthesis, than you would have, It should be C. x+7=7 both equations = 0

OpenStudy (anonymous):

so if you have, 5+x+2=7 x=0 C. x+7=7 x=0

OpenStudy (mathstudent55):

No one mentioned x = 0. The only question is which choice represents an equation that is equivalent to the given one. The goal is not to solve the equation; it's simply to choose what the next step in solving the equation is.

OpenStudy (anonymous):

Still, the question was, Which of the following equations is EQUIVALENT to (5 + x) + 2 = 7? EQUIVALENT referring to which equation is in the long run EQUAL to the given equation. The answer is still C.

OpenStudy (anonymous):

Actually, the fact that x=0 comes simply from solving the first equation. It doesn't matter what specific numbers and variables go into the equation, as long as they are equal when simplified a very simple equation and an extremely complicated one can be equivalent. So yes, the answer is C.

OpenStudy (anonymous):

Don't get me wrong I get what your saying and if I solved it, it was with the mare purpose to help him see how C was equivalent to the given equation.

OpenStudy (anonymous):

Thank you maud101 ^.^

OpenStudy (anonymous):

For sure